OpenStudy (anonymous):

I need help with calculus 2 integrals

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ sech(\ln(x))\tanh(\ln(x)) }{ x } dx\]

OpenStudy (perl):

did you try u = ln x

OpenStudy (anonymous):

du = 1/x dx

OpenStudy (anonymous):

What do I do with the sech and tanh though or do I just leave them there?

OpenStudy (perl):

now look for the integral of sech tanh

OpenStudy (perl):

what is the derivative of sech ?

OpenStudy (anonymous):

-sechxtanhx

OpenStudy (perl):

so use that

OpenStudy (perl):